The smaller of two consecutive even integers is five more than one-half of the greater. Find the integers
Answer by midwood_trail(310)

The smaller of two consecutive even integers is five more than one-half of the greater. Find the integers.
Let x = the smaller even integer
Let x + 2 = the larger even integer
The statement "The smaller of two consecutive even integers is five more" can be written simply as x + 5.
The statement "one-half of the greater" is written (1/2)(x + 2).
We now equate the two quantities and solve for x.
x + 5 = (1/2)(x + 2)
Can you take it from here?
When you find x, plug the value of x into the first and second even integers as written above.